Let f be a nonnegative function in L^1(R^n). Prove that
|f̂(ξ)| ≤ f̂(0), ξ∈R^n
1
Expert's answer
2019-10-04T10:33:27-0400

"|f̂(ξ)|=|\int_{-\infty}^{\infty}e^{-2\pi i x ξ}f(x)dx|\le"

"\int_{-\infty}^{\infty}|e^{-2\pi i x ξ}||f(x)|dx="

"\int_{-\infty}^{\infty}1\cdot f(x)dx="

"\int_{-\infty}^{\infty}|e^{-2\pi i x \cdot0}|f(x)dx="  f̂(0)

And therefore |f̂(ξ)| ≤ f̂(0).